Two masses m1=1kg and m2 =2kg are connected by suspended by a weightless string by a massless pulley .assume that both the masses start from rest the distance travelled by the centre of mass in 2s is -a) 20/9m b) 40/9m c) 2/3 m d) 1/3 m

Grade:12th

1 Answers

siddharth swami
6 years ago
mm2g- T = ma
T- m1g = m1a
solving these equations we get, a= g/3
distance travelled by each block
s= ut+ (1/2)at2 = 20/3 m
apllying centre of mass formula
(m2*s – m1*s)/(m2+m1)= 20/9